OpenStudy (anonymous):

Write the slope-intercept form of the line which passes through (4, 2) with a slope of 1/2.

OpenStudy (saifoo.khan):

\[y - 2 = \frac 12 (x-4)\]try solving..

OpenStudy (anonymous):

can u give me the answer? Because I have to go in a second.

OpenStudy (saifoo.khan):

y = x/2

OpenStudy (anonymous):

Thanks!

OpenStudy (anonymous):

You're a life saver! :)

OpenStudy (saifoo.khan):

Thank you!! =)

OpenStudy (anonymous):

:)